OpenStudy (anonymous):

What's the interest on $175.00 at 4% for 120 days?

OpenStudy (anonymous):

p=175.00 rate=4% time=120 days so Simpleinterest=p*r*t/100 and convert 120 days into years

OpenStudy (amistre64):

doesnt simple interest use a 360 day year?

OpenStudy (anonymous):

That's what I thought. But the problem is in days.

OpenStudy (amistre64):

120/360 = 1/3; so best guess is to use the time variable as 1/3